Marlin Motors sells a single product with a selling price of $490 with variable costs per unit of $196. The company's monthly fixed expenses are $52,920.

C. Prepare a contribution margin income statement for the month of November when they will sell 140 units. Use a minus sign for a net loss if present.

Income Statement
Sales $
Variable Costs
Contribution Margin $
Fixed Costs
Net Loss $

D. How many units will Marlin need to sell in order to realize a target profit of $61,740?

New break-even units  units

E. What dollar sales will Marlin need to generate in order to realize a target profit of $61,740?

New break-even dollars $

F. Construct a contribution margin income statement for the month of February that reflects $259,700 in sales revenue for Marlin Motors.

Income Statement
Sales $
Variable Costs
Contribution Margin $
Fixed Costs
Net Income $
Answer #1

C. Prepare a contribution margin income statement for the month of November when they will sell 140 units. Use a minus sign for a net loss if present.

Income Statement
Sales (140*490) $68600
Variable Costs (140*196) 27440
Contribution Margin $41160
Fixed Costs 52920
Net Loss -11760

D. How many units will Marlin need to sell in order to realize a target profit of $61,740?

New break-even units = (52920+61740)/294 = 390 units

E. What dollar sales will Marlin need to generate in order to realize a target profit of $61,740?

New break-even dollars $390*490 = $191100

F. Construct a contribution margin income statement for the month of February that reflects $259,700 in sales revenue for Marlin Motors.

Income Statement
Sales $259700
Variable Costs 103880
Contribution Margin $155820
Fixed Costs 52920
Net Income $102900